OpenStudy (anonymous):

a bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probablitly that you pick a red marble, do not replace it, and pick another red marble?

OpenStudy (kirbykirby):

1st marble picked: There are 4+3+3 =10 marbles total. The prob. of getting one red is 3/10. 2nd marble: There are now 9 marbles left. Now one red was picked before, so now the probability of getting a red is 2/9. Now you use the product rule since you want "red on 1st pick AND red on 2nd pick" = 3/10*2/9 = 1/15